# DynamicPAE: Generating Scene-Aware Physical Adversarial Examples in Real-Time

### v1

@@ 3.1 Residual-Driven Sample Trajectory Guidance
where $\gamma$ is the hyper-parameter. Note that the latent regularization is more like BatchNorm, which aims at tackling the optimization difficulties rather than preventing overfitting. We further compare the regularization with the classical VAE prior in experiments.
3.2 Context-Aligned Scene Expectation Simulation
The practical PAE methods should be compatible with the variable and agnostic scenario. For the dynamic attack scenarios, the critical issues are incomplete observations from the attacker, which limit the perception of the generator, and the agnostic nature of target models and tasks, which affect the balance between objectives. To bridge the gap between the training environment and potential deployment scenarios, firstly, we propose the conditional-uncertainty-aligned data module. This module allows the generator to accommodate incomplete observations by employing a conditional probabilistic model for domain randomization. Secondly, we propose the skewness-aligned objective re-weighting module, which uses the skewness indicator to balance multiple optimization objectives, including stealth and aggressive, under agnostic attack targets. Thus, consistency in the sampling and control of stealth and attack objectives is maintained, and hyper-parameter tuning is eliminated.
@@ 3.2 Context-Aligned Scene Expectation Simulation
$\bullet$ Conditional-Uncertainty-Aligned Data Model: We align the training and real-world scenes by creating a conditional probabilistic model to generate the incomplete observation $\mathbf{P}_{\mathrm{X}}$, as illustrated in Figure 3, based on the dynamic attack problem formulation in Equation 2. In this model, the parameters $\boldsymbol{\theta}$ of the attack injection process are generated by the current state of the world $\mathbf{X}$ and a random factor $\boldsymbol{s}$, which, along with $\mathbf{X}$ and another random factor $\boldsymbol{s}^{\prime}$, jointly generate the physical context data $\mathbf{P}_{\mathrm{X}}$ that the attacker acquires. The training is performed by minimizing the expected losses under the data model, where $\Omega$ and $\Theta$ are the probability distribution that generates $\mathbf{P}_{\mathrm{X}}$ and $\boldsymbol{\theta}$, respectively:
@@
$\mathbf{X}\oplus(\boldsymbol{\delta},\boldsymbol{\theta}):=$ $\mathbf{X}\otimes(1-\boldsymbol{m}_{\theta})+$ (13)
$\textit{AffineTransform}\left(\boldsymbol{\delta},A_{\boldsymbol{\theta}}\right)\otimes\boldsymbol{m}_{\theta},$
@@
where $\boldsymbol{s}^{\prime}$ is the generated random seed for RandAugment.$\Theta$ is constructed by sampling the object location according to the objects in $\mathbf{X}$, and $\mathbf{X}$ is constructed by the training dataset of the target model.
$\bullet$ Skewness-Aligned Objective Re-weighting: The dynamic PAE generator may be deployed to evaluate and attack diverse victim models and corresponding tasks. To balance the stealth and attack objectives, achieve controllable stealthiness, and avoid manual adjustment of the optimization-related parameters when facing new targets, we further propose an objective re-weighting approach based on the skewness measurement [55]. Specifically, we first focus on adding and determining the additional relative weights $\alpha\in(0,1)$ of the residual task, and define the optimization objective using the residual fusion operator in Eq. 6 as: $\mathscr{L}_{\lambda}((1-\alpha)\mathcal{L}_{\mathrm{Atk}},\alpha\mathcal{L}_{\mathrm{Inv}})=\boldsymbol{\lambda}(1-\alpha)\mathcal{L}_{\mathrm{Atk}}+(1-\boldsymbol{\lambda})\alpha\mathcal{L}_{\mathrm{Inv}}$.
@@
Figure 4: Skewness-based Hyperparemeter Control
Intuitively, skewness aims to measure the lack of symmetry in the distribution of a dataset around its mean. On one hand, as shown in Figure 4a, if the skewness is positive, the distribution is right-skewed or positively skewed, and vice versa. On the other hand, as shown in Figure 4b, the $\mathcal{L}_{\mathrm{Inv}}$ of the samples that are generated by middle-size $\lambda$ tends to decrease within the optimal curve as the weights $\alpha$ increase, leading to a decrease in the loss skewness. This allows for the skewness to be utilized as an indicator variable to control the value of $\alpha$ by measuring the distribution of losses. Thus, based on the skewness of the loss values, we define the gradient feedback of the $\alpha$-controller of loss $\leavevmode\nobreak\ \mathcal{L}$ as:
$\nabla\alpha$ $:=-\text{skewness}(\mathcal{L}_{\mathrm{Inv}})+\varsigma,$ (15)
$\text{skewness}(\mathcal{L_{\mathrm{Inv}}})$ $:=\mathbb{E}[\frac{(\mathcal{L}_{\mathrm{Inv}}-\mu(\mathcal{L}_{\mathrm{Inv},\lambda=0,1}))^{3}}{\sigma(\mathcal{L}_{\mathrm{Inv},\lambda=0,1})^{3}}],$
where $\varsigma$ is the target skewness, $\mu$ and $\sigma$ are the mean and standard deviation of the loss when $\lambda\in\{0,1\}$, which replaces the statistics for all samples for better stability. The gradient $\nabla\alpha$ equals zero if and only if the loss is balanced, that is, $\text{skewness}(\mathcal{L})=\varsigma$. For implementation, we re-parameterize $\alpha$ as $\alpha=e^{\beta}\in(0,+\infty)$ and optimize $\beta$ using Adam optimizer. We set the target skewness of the reconstruction loss $\mathcal{L}_{\mathrm{Inv}}$ to zero ($\varsigma=0$), and the estimation of $\mu$ and $\sigma$ is performed by the moving average. The convergence of the controller is shown by the following proposition, and the proof is given in the supplementary material by analyzing the trade-off curve.
@@
$\nabla^{2}\alpha<0$ if $\mathcal{L}_{\mathrm{Atk}}$ is strictly negatively correlated to $\mathcal{L}_{\mathrm{Inv}}$ for any set of measured PAEs $\boldsymbol{\delta}_{\lambda}$ that are optimal for the overall optimization objective $\mathcal{L}$ under a certain value of $\lambda$.
We further generalize our framework to the multi-objective attack scenario, consisting of $t$ attack and residual tasks. The losses of these tasks are composed by the sample-wise loss weights $\boldsymbol{\lambda}=[\boldsymbol{\lambda}_{1},\boldsymbol{\lambda}_{2},...,\boldsymbol{\lambda}_{b}]\in\mathbb{R}^{b\times t}$, which are encoded into $\mathcal{G}$, and task-wise loss weights $\boldsymbol{\alpha}\in(\mathbb{R^{+}})^{t}$, which are controlled by the skewness indicator and satisfy $\sum_{i=1}^{t}\alpha_{i}=1$. We define the composed loss as the Guided and Controlled Attack Loss $\mathcal{L}_{\mathrm{GCAtk}}$:
$\mathcal{L}_{\mathrm{GCAtk}}:=\sum_{i=1}^{b}\sum_{j=1}^{t}\{\boldsymbol{\lambda}\boldsymbol{\alpha}^{T}\otimes\mathbf{L}\}_{i,j},$ (16)
where $\otimes$ denotes the Hadamard product, $t$ is the number of tasks, $b$ is the batch size, and $\mathbf{L}\in\mathbb{R}^{b\times t}$ is the loss matrix of $t$ tasks and $b$ items in a batch defined by the attack scenario with $\mathcal{L}_{\mathrm{Atk}}$. For the single-objective adversarial attack implementation, $t=2$ and $\mathbf{L}=[\mathcal{L}_{\mathrm{Atk}},\mathcal{L}_{\mathrm{Inv}}]$, and in this case:
@@
$\mathcal{L}_{\mathrm{GCAtk}}=\mathscr{L}_{\lambda}((1-\alpha)\mathcal{L}_{\mathrm{Atk}},\alpha\mathcal{L}_{\mathrm{Inv}}).$ (17)
For the attack loss $\mathcal{L}_{\mathrm{Atk}}$, we implement it according to the specific scenario in experiments.
$\bullet$ Training and Deployment Pipeline: By aligning the training environment and deployment environment, we can establish an end-to-end pipeline for building dynamic PAE generator. For the overall training, the total loss is formulated as:
$\mathcal{L}_{\mathrm{Total}}=\mathcal{L}_{\mathrm{GCAtk}}+\mathcal{L}_{\mathrm{Reg}}+\mathcal{L}_{\mathrm{VAE}},$ (18)
where $\mathcal{L}_{\mathrm{Reg}}$ is the regularization loss in Eq. 11. We also incorporate an auto-regressive training loss $\mathcal{L}_{\mathrm{VAE}}$ with $\beta$-VAE [56] with images from the dataset $\mathcal{X}$ to train $\mathrm{Enc}$ and let $\mathcal{G}$ maintain the modeling capability of more complex natural images. The training pipeline is performed by simulating the physical observation $\mathbf{P}_{\mathrm{X}}$ and sampling $\boldsymbol{\lambda}$, generating PAEs $\boldsymbol{\delta}$ based on it, evaluating the losses, and performing optimization on model parameters by $\nabla\mathcal{L}_{\mathrm{Total}}$ and on $\boldsymbol{\alpha}$ by Eq. 15. We deploy the generator directly on edge devices after training and achieve real-time generation by connecting it with the camera and display. Detailed training implementation is shown in the supplementary material.
@@ 4 Experiments
We evaluate our framework in both digital and physical environments on the object detection task since it is classical for both physical and digital attacks. The experiments are conducted in the following aspects: ❶ The performance of our model in the benchmark. ❷ The analysis corresponds with the methodologies that tackle the challenge of noisy gradient feedback in PAE generator training and the agnostic nature of attack scenarios. ❸ The ablation of the key components.
@@ 4.1 Experiment Settings
In this subsection, we describe the key settings of the experiments. We use the COCO [57] and the Inria [58] as the dataset of the person detection task, while data collected from CARLA is used for the simulation experiment. Detailed attack transformation, method implementation, dataset description, and victim model description are shown in the supplementary material.
Figure 5: Aggressiveness-distortion curve evaluated on COCO dataset, corresponding to Table III.
Metrics: The average precision (AP), formulated as follows by the precision $Pre(c)$ and recall $Rec(c)$ of confidence threshold $c$ is applied as the metric of adversary performance.
@@
We adopt two confidence thresholds: the first is $c_{min}=50\%$, with the corresponding metric denoted as $AP_{50}$, which serves as the evaluation metric identical to the practice of previous object detection attack evaluation. The second threshold is $c_{min}=1\%$, with the metric denoted as $AP_{01}$, as a more stringent evaluation metric. All $\boldsymbol{iou}$ thresholds are set as 0.5. The low-confidence and unattacked bounding boxes are filtered out during post-processing.
To ensure fairness, we use consistent patch placement parameters in the evaluation to control the overall stealthiness of the attack. Some specific methods also have other factors affecting stealthiness, such as the naturalness constraint of GAN-NAP, so we additionally assessed relevant metrics in the experiment. We use the Structural Similarity Index (SSIM) and the Learned Perceptual Image Patch Similarity (LPIPS) as metrics to measure the stealthiness of the patches. Both SSIM and LPIPS have a range of [0, 1]. SSIM is not included as a loss during the training of our model.
@@
Baselines PGD AdvGAN T-SEA GAN-NAP
DNN based
Physical Capability
@@
Baseline Attack Methods: For the novel task of dynamic attack, we select representative methods as baselines and adapt them to the test setup including PGD [59], AdvGAN [37], GAN-NAP [42], also known as NAP, and T-SEA [43]. The comparison of these methods is shown in Table II. PGD, GAN-NAP, and T-SEA are adapted as patch-form PAE evaluation baselines. For the simulated experiment, we conducted supplementary evaluations on similar AdvGAN methods. For the physical experiment, we performed tests using existing patch attacks (GAN-NAP and T-SEA).
Attack Method Implementations: Unlike training universal adversarial patch, dynamic PAE training requires the model to learn spatial correlations, thus a selection of bounding boxes (Bboxes) is needed. For loss construction, the maximum confidence value is employed following GAN-NAP with the addition of a selection mask:
$\mathcal{L}_{\mathrm{Atk}}=||\boldsymbol{conf}\cdot[\boldsymbol{iou}>0.3]||_{\infty},$ (20)
where $\boldsymbol{iou}$ is calculated between the generated and target Bboxes. The detailed implementation is provided in the supplementary material.
TABLE III: Performance Evaluation. indicates smaller values are better, while signifies larger values are preferred. Since the patch size is already constrained, stealthiness is only for better fairness.

### v3

@@ 3.1 Residual-Guided Adversarial Pattern Exploration
where $\gamma$ is the hyperparameter. Note that the latent regularization is more like BatchNorm, which aims at tackling the optimization difficulties rather than preventing overfitting. We further compare the regularization with the classical VAE prior in experiments.
3.2 Distribution-Matched Attack Scenario Alignment
Next, we focus on the alignment problem between $\mathcal{G}$ and the real-world adversarial attack scenario. Regarding the trade-off, the optimal attack performance is highly related to the universality and the stealthiness; the training environment of $\mathcal{G}$ shall align these properties with the requirements of the real-world scenario. Therefore, we regard the alignment problem as the unbiased construction problem of the training data and objectives. To solve the problem, we propose the Distribution-Matched Attack Scenario Alignment as a design methodology, consisting of the conditional uncertainty alignment and the objective bias alignment. The former controls the attack universality by aligning the sampled scenes’ distribution and the real-world environment, while the latter controls the attack stealthiness by aligning the skewness statistics of the generator’s loss distribution.
@@ 3.2 Distribution-Matched Attack Scenario Alignment
$\bullet$ Conditional-Uncertainty-Aligned Data Model: Bringing EoT [23] and domain randomization [25], we align the training and real-world scenes by creating a conditional probabilistic model to generate both of the parameters of the attack injection process $\boldsymbol{\theta}$ and the attackers’ observation $\mathbf{P}_{\mathrm{X}}$, as illustrated in Figure 4, based on the problem formulation in Equation 2. In this model, $\boldsymbol{\theta}$ are generated by the current state of the world $\mathbf{X}$ and a random factor $\boldsymbol{s}$, which, along with $\mathbf{X}$ and another random factor $\boldsymbol{s}^{\prime}$, jointly generate the physical context data $\mathbf{P}_{\mathrm{X}}$ that the attacker acquires. The training is performed by minimizing the expected losses under the data model, where $\Omega$ and $\Theta$ are the probability distribution that generates $\mathbf{P}_{\mathrm{X}}$ and $\boldsymbol{\theta}$, respectively:
@@
$\mathbf{X}\oplus(\boldsymbol{\delta},\boldsymbol{\theta})$= $\mathbf{X}\otimes(1-\boldsymbol{m}_{\theta})+$ (13)
$\textit{AffineTransform}\left(\boldsymbol{\delta},A_{\boldsymbol{\theta}}\right)\otimes\boldsymbol{m}_{\theta},\$
@@
[…]
$\bullet$ Skewness-Aligned Objective Re-weighting: The dynamic PAE generator may be deployed to evaluate and attack diverse victim models and corresponding tasks. To balance the stealth and attack objectives in PAE exploration, achieve controllable stealthiness, and avoid manual adjustment of the optimization-related parameters when facing new targets, we further propose an objective re-weighting approach based on the skewness measurement [48]. Specifically, we focus on adding and determining the additional relative weights $\alpha\in(0,+\infty)$ of the residual task, and define the optimization objective using the residual fusion operator in Eq. 6 as: $\mathscr{L}_{\lambda}(\mathcal{L}_{\mathrm{Atk}},\alpha\mathcal{L}_{\mathrm{Inv}})=\boldsymbol{\lambda}\mathcal{L}_{\mathrm{Atk}}+(1-\boldsymbol{\lambda})\alpha\mathcal{L}_{\mathrm{Inv}}$.
@@
Figure 5: Skewness-based Hyperparameter Control
[…]
$\nabla\alpha$ $=-\text{skewness}(\mathcal{L}_{\mathrm{Inv}})+\varsigma,$ (15)
$\text{skewness}(\mathcal{L_{\mathrm{Inv}}})$ $=\mathbb{E}[\frac{(\mathcal{L}_{\mathrm{Inv}}-\mu(\mathcal{L}_{\mathrm{Inv},\lambda=0,1}))^{3}}{\sigma(\mathcal{L}_{\mathrm{Inv},\lambda=0,1})^{3}}],$
where $\varsigma$ is the target skewness, $\mu$ and $\sigma$ are the mean and standard deviation of the loss when $\lambda\in\{0,1\}$, which replaces the statistics for all samples for better stability. The gradient $\nabla\alpha$ equals zero if and only if the loss is balanced, that is, $\text{skewness}(\mathcal{L})=\varsigma$. We apply the gradient descent with gradient defined by $\nabla\alpha$ to optimize $\alpha$, resulting in a closed-loop control of loss distribution on explored PAEs $\boldsymbol{\delta}\sim\mathcal{G}(\cdot)$, as shown in Figure 6. The following proposition shows the convergence of the optimization. Detailed assumptions and proof are provided in the supplementary material through an analysis of the trade-off curve.
Figure 6: Closed-loop Control of Loss Distribution
@@
$\nabla^{2}\alpha>0$ if $\mathcal{L}_{\mathrm{Atk}}$ is strictly negatively correlated to $\mathcal{L}_{\mathrm{Inv}}$ for any set of measured PAEs $\boldsymbol{\delta}_{\lambda}$ that are optimal for the overall optimization objective $\mathcal{L}$ under a certain value of $\lambda$.
For implementation, we set the target skewness of the reconstruction loss $\mathcal{L}_{\mathrm{Inv}}$ to zero ($\varsigma=0$). We use the shifted window algorithm to measure $\mu(\cdot)$ and $\sigma(\cdot)$, and then the skewness statistics, i.e, the expectation $\mathbb{E}$ is estimated from the loss of recent trained batches. Then, we use the Adam optimizer with learning rate lr=1e-3 to optimize $\alpha$, and reparameterize $\alpha$ as $\alpha=e^{\beta}\in(0,+\infty)$ for the domain constraint. An equivalent loss is constructed to perform optimization along with $\mathcal{G}$ at each training iteration:
$\textit{Adam}[{\alpha}]$ $\text{.zero\_grad()};$ (16)
$\mathcal{L}(\alpha)$ $=\mathrm{float}(\nabla\alpha)*\alpha;$
$\mathcal{L}(\alpha)$ $\mathrm{.backward()};$
$\textit{Adam}[{\alpha}]$ $\text{.step()}.$
3.3 Neural Architecture and Training Implementations
Based on the introduced training technology, we realize the ongoing adaptation of the typical patch-based PAE generation through a lightweight construction of neural architecture and a modularized construction of the training pipeline. After training, we deploy the generator and enable the real-time generation by connecting it to a camera and display. Details are shown in the supplementary material.
$\bullet$ Neural Architecture: For the neural generator $\mathcal{G}_{\boldsymbol{\lambda}}(\mathbf{P}_{\mathrm{X}})$ $\mathrm{Dec}(\mathrm{Enc}_{\mathbf{P}}(\mathbf{P}_{\mathrm{X}},\boldsymbol{\lambda}))$, we constructed a lightweight physical environment encoder $\mathrm{Enc}_{\mathbf{P}}$ based on the multi-modal information fusion paradigm. Specifically, the encoding model of $\mathbf{P}_{\mathrm{X}}=(\mathbf{X}_{local}^{\prime},\mathbf{X}_{global}^{\prime},\boldsymbol{m}^{\prime})$ is constructed by:
$\mathrm{Enc}_{\mathbf{P}}=\textit{MLP}_{1}(\textit{LN}($ $\textit{MLP}_{2}(\textit{ResNet}_{1}(\mathbf{X}_{local}^{\prime})))$ (17)
$+\textit{LN}($ $\textit{MLP}_{3}(\textit{ResNet}_{2}(\text{Concat}[\mathbf{X}_{global}^{\prime},\boldsymbol{m}^{\prime}]))$
[…]
where Resnet is for feature extraction, MLP denotes the multi-layer perception module that models the mapping, and LN denotes the layer-norm module that regularizes the feature importance. $\mathrm{Dec}$ is a lightweight version of BigGAN [49] decoder, with the self-attention module removed.
$\bullet$ Training Pipeline: Following the practice of multi-task fine-tuning, we train the generator $\mathcal{G}$ with the VAE-pretraining task and the new residual-guided adversarial attack task. The total loss is formulated as:
$\mathcal{L}_{\mathrm{Total}}=\mathscr{L}_{\lambda}(\mathcal{L}_{\mathrm{Atk}},\alpha\mathcal{L}_{\mathrm{Inv}})+\mathcal{L}_{\mathrm{Reg}}+\mathcal{L}_{\mathrm{VAE}},$ (18)
where $\mathscr{L}_{\lambda}(\cdot):=\lambda\mathcal{L}_{\mathrm{Atk}}+(1-{\lambda})\alpha\mathcal{L}_{\mathrm{Inv}}$ is the residual-guided training loss, $\mathcal{L}_{\mathrm{Reg}}$ is the regularization in Eq. 11. $\mathcal{L}_{\mathrm{VAE}}$ is generative training loss of $\beta$-VAE [50], and we use it to train $\mathrm{Dec}$ with images from the dataset $\mathcal{X}$. It lets $\mathcal{G}_{\lambda}$ maintain the modeling capability of more complex natural images and improves the naturalness of the generated PAEs when the attack loss weight $\lambda$ is low. The overall training is shown in Algorithm 1. The modularized implementation of the training and deployment pipeline is shown in Figure 7. The gray modules represent existing components in the deep learning and adversarial attack pipeline, while the purple modules are the new components to be constructed. The pipeline is straightforward, as no additional discriminators have been introduced. $\mathcal{L}_{\mathrm{VAE}}$ is not shown since it is optional.
Figure 7: Training and deployment pipeline. The arrow in the training pipeline denotes the forward process in a training step, and the gradient optimization is performed on $\mathcal{G}$. The deployment pipeline processes the data and displays PAE as a real-time stream.
1
Input: Dataset $\mathcal{D}\subset\textbf{Images}\times\textbf{BBoxes}$, Target model $\mathcal{F}$ with attack loss $\mathcal{L}_{\mathrm{Atk}}$, Untrained PAE generator $\mathcal{G}$
Output: Trained model $\mathcal{G}$
2 $\alpha\leftarrow 1.0$
3for $epoch\in\{1,2...n\}$, $step\in\{1,2...m\}$ do
4   Sample $X$ from the dataset $\mathcal{D}$.
5  Sample the patch location $\boldsymbol{\theta}$}.
6  Sample task ratios $\boldsymbol{\lambda}$ by Equation 9.
7  Pre-process $\mathbf{P}_{\mathrm{X}}$ by Equation 14.
8  Generate PAEs $\boldsymbol{\delta}_{\boldsymbol{\lambda}}\leftarrow\mathrm{Dec}\circ\mathrm{Enc}_{\mathbf{P}}(\mathbf{P}_{\mathrm{X}},\boldsymbol{\lambda})$.
9 $\mathbf{X}^{\prime}\leftarrow\mathbf{X}\oplus(\boldsymbol{\delta}_{\boldsymbol{\lambda}},\boldsymbol{\theta})$. (Equation 13)
10 $\mathcal{L}_{\mathrm{Inv}}\leftarrow\mathrm{MSE}(\mathbf{X},\mathbf{X}^{\prime})+\mathrm{LPIPS}(\mathbf{X},\mathbf{X}^{\prime})$.
@@ 3.3 Neural Architecture and Training Implementations
11 $\mathscr{L}_{\boldsymbol{\lambda}}\leftarrow\boldsymbol{\lambda}\mathcal{L}_{\mathrm{Atk}}(\cdot)+(1-\boldsymbol{\lambda})\alpha\mathcal{L}_{\mathrm{Inv}}$.
[…]
13 $\mathcal{L}_{\mathrm{VAE}}\leftarrow\beta$-VAE loss [50] on $\mathrm{Dec}$ with data $\mathbf{X}$.
14 $\mathcal{L}_{\mathrm{Total}}\leftarrow\mathscr{L}_{\boldsymbol{\lambda}}+\mathcal{L}_{\mathrm{Reg}}+\mathcal{L}_{\mathrm{VAE}}$.
15  Optimize $\mathcal{G}$ with $\mathcal{L}_{\mathrm{Total}}$ using Adam optimizer.
16  Optimize $\alpha$ with $\mathcal{L}_{\mathrm{Inv}}$ by Equation 15, 16.
17return $\mathcal{G}$.
Algorithm 1 Ends-to-Ends DynamicPAE Training
@@ 4 Experiments
We evaluate our framework in both digital and physical environments on the object detection task since it is classical for both physical and digital attacks. The experiments are conducted in the following aspects: ❶ The performance of our model in the benchmark. ❷ The analysis corresponds with the methodologies that tackle the challenge of noisy gradient feedback in training and the alignment issues between trained generators and attack scenarios. ❸ The ablation of the key components.
@@ 4.1 Experiment Settings
In this subsection, we describe the key settings of the experiments. We use the COCO [51] and the Inria [52] as the dataset of the person detection task, while data collected from CARLA is used for the simulation experiment. Detailed method implementation, dataset description, and victim model description are shown in the supplementary material.
Metrics: The average precision (AP), formulated as follows by the precision $Pre(c)$ and recall $Rec(c)$ under the confidence threshold $c$, and average success rate (ASR) are applied as the metric of attack performance.
@@
We adopt two confidence thresholds: the first is $c_{min}=50\%$, with the corresponding metric denoted as $AP_{50}$, which serves as the evaluation metric identical to the practice of previous object detection attack evaluation. The second threshold is $c_{min}=1\%$, with the metric denoted as $AP_{01}$, as a more stringent evaluation metric. All $\boldsymbol{iou}$ thresholds are set as 0.5. The low-confidence and unattacked bounding boxes are filtered out during post-processing. For ASR, the attack is regarded as successful if the target’s detection confidence is below 0.5.
We use the Structural Similarity Index (SSIM) and the Learned Perceptual Image Patch Similarity (LPIPS) as metrics to measure the stealthiness of the patches. Both SSIM and LPIPS have a range of [0, 1]. SSIM is not included as a loss during the training of our model.
@@
Baselines PGD AdvGAN GAN-NAP T-SEA CAP
Optimization Space Image CNN Latents Image Image*
@@
Baseline Attack Methods: Our baselines include PGD [53], AdvGAN [31], GAN-NAP [37], also known as NAP, T-SEA [38] and CAP [54]. Each of them represents a distinct optimization method. The comparison is shown in Table II, where CAP applies a min-max optimization together with the ISP surrogate network. For the strict physical patch attack scenario that matches our design, GAN-NAP, T-SEA, and CAP are included. Although not physically applicable due to computational issues, we also include PGD (with single-image EoT [23]), as there is no prior adaptive physical adversarial patch. To further validate our technical contribution, we supplement additional comparison with GDPA [55] on real-time digital adversarial patch generation for face recognition in Section IX, and with AdvGAN on real-time perturbation generation in Section 4.6.
Attack Method Implementations: For loss construction of DynamicPAE  the maximum confidence value is employed following GAN-NAP with the addition of a selection mask. We select the bounding box by the $\boldsymbol{iou}$ between the detector’s generated bbox and the target person’s bbox:
[…]
Detailed implementations of DynamicPAE and baselines are in the supplementary material.
TABLE III: Spatial Transformation Settings. Position and size are related to the annotated bounding-box.
Name Position Size Rotation 3D Rotation
Base [0.4, 0.6] [0.04, 0.06] [-9°, +9°] 0
S+ [0.4, 0.6] [0.03, 0.07] [-9°, +9°] 0
P+ [0.3, 0.7] [0.04, 0.06] [-9°, +9°] 0
AF [0.4, 0.6] [0.04, 0.06] [-9°, +9°] [-30°, +30°]
Zero 0.5 0.05 0 0
TABLE IV: Color Transformation Settings
Brightness Saturation Contrast Gaussian Noise
Param. ±0.05 ±0.1 ±0.1 [0, 0.02]
Attack Simulation: To ensure fairness, we use consistent patch placement parameters in the evaluation. The patch placement model for evaluation is based on affine transformation $A$, which can be decomposed into rotation matrices, scaling matrices, and translation matrices:
$A=A_{rotate}\cdot A_{scale}\cdot A_{trans}.$ (21)
Thus, we define separate parameter distribution models for rotation, scaling, and translation in the transformation. Specifically, we sample the relative position, the relative size, the rotation in the 2D image, and the rotation in the 3D space (the rotation axis is parallel to the camera’s imaging plane), from a predefined uniform distribution shown in Table III. Without specification, Base transformation is applied in the following experiments. In comparison to previous evaluations, a smaller patch size was implemented across the evaluated attacks, which increases the attack difficulty. For color distortion, we apply brightness, saturation, and contrast transformation with uniform sampled parameters, and then we add Gaussian noise with standard variance sampled from the uniform distribution, as shown in Table IV.
TABLE V: Performance Evaluation. indicates smaller values are better, while signifies larger values are preferred. Since the patch size is already constrained, stealthiness is only for better fairness.